OpenStudy (anonymous):

Calculus Fan and Medal! If e^(xy) = 4, then what is dy/dx at the point (1, ln4)?

OpenStudy (welshfella):

this is tricky let u = xy by chain rule du /dx = d e^u * du ---- -- dx dx

OpenStudy (welshfella):

* thats d e^u / du

OpenStudy (welshfella):

e ^xy * d (xy) = d (4)/dx dx

OpenStudy (anonymous):

Okay so I basically solve for the implicit derivative of that equality? From there how do I bring in the points given?

OpenStudy (welshfella):

just plug in x = 1 and y = ln 4

OpenStudy (welshfella):

d(xy)/dx = y + x dy/dx ( product rule)